OpenStudy (anonymous):

The cost of printing a store flyer is $450 for 1000 copies and $880 for 2000 copies. Write a linear equation modeling this data. Then, find the amount of money it would cost to print 6245 copies.

OpenStudy (anonymous):

450 = 1000m + b 880 = 2000m + b Subtracting 430 = 1000m OR m = 0.430 and b = 450 -- 1000*0.430 = 20 Equation for n copies is C(n) = $[0.430n + 20] C(6245) = $[6245*0.430 + 20] = $2705.35

OpenStudy (whpalmer4):

Think of (copies, price) as (x, y). Now find the slope and y-interct from those two points, write the new equation, and use it to find out the cost of the big print job.